OpenStudy (anonymous):

write a mathematical expression for each phrase and simplify the expression. 1) the sum of eight and the quotient of ten and five 2) the sum of one-half and the reciprocal of two-thirds

OpenStudy (anonymous):

1) the sum of eight and the quotient of ten and five 10 8 + ---- = 8 + 2 = 10 5 2) the sum of one-half and the reciprocal of two-thirds 1 2 3 one half = --- and reciprocal of --- = ----- 2 3 2 1 3 1 + 3 4 --- + --- = ------ = ---- = 2 2 2 2 2